What is the derivative of y=(x + tanx)÷(secx + tanx)

Grade:11

1 Answers

Shivansh Saumya
7 years ago
y = (x+tanx)/(secx+tanx)
dy/dx = d((x+tanx)/(secx+tanx))/dx
quotient rule for derivatives = (f(x)/g(x))’ = g(x).f’(x) – f(x).g’(x)/(g(x))2
= ((secx+tanx).d(x+tanx)/dx – (x+tanx).d(secx+tanx)/dx)/(secx+tanx)2
=((secx+tanx).(1+sec2x) – (x+tanx).(secxtanx+sec2x))/(secx+tanx)2
solve it to get the answer. also refer to arihant aieee